# Code-in-the-Loop Forensics: Agentic Tool Use for Image Forgery Detection

### v1

@@ Abstract
Existing image forgery detection (IFD) methods either exploit low-level, semantics-agnostic artifacts or rely on multimodal large language models (MLLMs) with high-level semantic knowledge. Although naturally complementary, these two information streams are highly heterogeneous in both paradigm and reasoning, making it difficult for existing methods to unify them or effectively model their cross-level interactions. To address this gap, we propose ForenAgent, a multi-round interactive IFD framework that enables MLLMs to autonomously generate, execute, and iteratively refine Python-based low-level tools around the detection objective, thereby achieving more flexible and interpretable forgery analysis. ForenAgent follows a two-stage training pipeline combining Cold Start and Reinforcement Fine-Tuning to enhance its tool interaction capability and reasoning adaptability progressively. Inspired by human reasoning, we design a dynamic reasoning loop comprising global perception, local focusing, iterative probing, and holistic adjudication, and instantiate it as both a data-sampling strategy and a task-aligned process reward. For systematic training and evaluation, we construct FABench, a heterogeneous, high-quality agent-forensics dataset comprising 100k images and approximately 200k agent–interaction question-answer pairs. Experiments show that ForenAgent exhibits emergent tool-use competence and reflective reasoning on challenging IFD tasks when assisted by low-level tools, charting a promising route toward general-purpose IFD. The code will be released after the review process is completed.
@@ 1 Introduction
In this work, we present ForenAgent, a novel interactive multi-turn framework that empowers MLLM to autonomously generate, execute, and iteratively refine Python-based low-level tools for IFD. To achieve this, we first abstract and generalize several commonly used low-level Python tools for IFD, such as frequency residual, noise residual, and high-pass filtering, and then integrate them into a comprehensive toolbox consisting of 12 candidate utilities for future community extension. As illustrated in Figure 1, ForenAgent autonomously orchestrates Python tools to verify a forged image from global screening to local inspection, ultimately classifying it as tampered and accurately localizing the forged region. The agent further demonstrates reflective self-correction by recovering from an initially misfocused crop to the correct area of interest, an “aha moment” observed in IFD agents.
The development of ForenAgent involves two key components: (1) Forgery Agent Benchmark (FABench), a high-quality and heterogeneous forensic agent dataset constructed using state-of-the-art generative models (e.g., GPT-4o [jaech2024openai], Nano-Banana [nano], and Midjourney-v7 [midjourney2023])). It contains 100k images (40k real, 30k synthetic, and 30k tampered) and serves as a comprehensive benchmark for training and evaluation in IFD. (2) A Cold-Start and Reinforcement Fine-Tuning (RFT) framework, designed to train MLLMs to function as reliable and autonomous agents. During the Cold-Start stage, ForenAgent adopts a self-exploration and experience-distillation paradigm. Specifically, GPT-4.1 [openai2025gpt41] observes a large collection of forgery samples from FABench under system prompts that provide procedural guidance and executable code examples, distilling operational patterns into structured agent–interaction training data for initialization. During RFT, we abstract the human IFD workflow into four reasoning stages: global perception, local focusing, iterative probing, and holistic adjudication. Correspondingly, we design four Forgery Process Rewards that together form the overall tool reward. By incorporating these verifiable reward components into the reinforcement learning process, ForenAgent develops a more interpretable and systematic forensic reasoning mechanism, effectively integrating basic image processing with low-level forensic analysis in a coherent investigative workflow. This design enables ForenAgent to explore diverse reasoning strategies and optimize for long-term process quality rather than simply imitating predefined answers.
@@
(1) We propose ForenAgent, a novel interactive, multi-turn framework that enables an MLLM to autonomously generate, execute, and iteratively refine Python-based low-level tools for image forgery detection, thereby taking the first step toward intelligent, tool-augmented IFD systems. (2) We construct FABench, a large-scale, high-quality, and heterogeneous forensic agent dataset comprising 100k images and 200k interactive QA pairs, focusing on forgery detection from cutting-edge generative models. (3) We formulate a dynamic reasoning loop comprising global perception, local focusing, iterative probing, and holistic adjudication into a data-sampling strategy and task-aligned process reward that foster flexible, tool-adaptive evidence reasoning and enable robust, interpretable decisions.
@@ 2.1 Image Forgery Detection
Early IFD research largely centered on low-level feature extractors that capture non-semantic inconsistencies between manipulated and authentic regions. Representative methods include frequency- and residual-based cues: FreqNet applies FFT to learn high-frequency patterns and improves cross-dataset generalization [tan2024frequency]; RGB-N employs Steganalysis Rich Model (SRM) filters to expose local noise inconsistencies [luo2021generalizing]; SAFE leverages Discrete Wavelet Transform (DWT)/Discrete Cosine Transform (DCT) with crop-based preprocessing to preserve local artifacts [luo2021generalizing]; PRNU-based approaches exploit camera fingerprints for multi-scale trace analysis with end-to-end fusion [zhang2023prnu]; JPEG ghost detection reveals regions with inconsistent compression [popescu2005exposing]; ObjectFormer localizes forgeries via DCT high-pass cues [Objectformerwang2022objectformer]; MVSS-Net integrates Bayar convolution and Sobel operators for fine-grained boundaries [MVSS-Net2021image], while HiFi-Net strengthens multi-level detection with high-pass filtering [HiFi_IFDL]. Additional work further enriches low-level forensic cues and robustness [chen2022self, bayar2016deep, liu2024forgery]. However, these methods often fail to capture higher-level forensic semantics, particularly those driven by language and knowledge, reducing their effectiveness against complex manipulations.
To address these challenges, recent work integrates large language models (LLMs) with vision–language reasoning for image forensics. For tampered images, Zhang et al. introduce DD-VQA [zhang2024common], a dataset of 2,968 FaceForensics++ samples used to fine-tune BLIP with contrastive loss, improving both accuracy and explanatory quality. FakeShield [xu2024fakeshield] combines an LLM with visual understanding via dual modules, while ForgeryGPT [liu2024forgerygpt] tailors an LLM to capture higher-order forensic cues across diverse feature spaces for explainable generation and interactive dialogue. For synthetic images, FakeScope [li2025fakescope] and LEGION [kang2025legion] leverage MLLMs to deliver strong explainability and detection performance. Building on these advances, SIDA [huang2025sidasocialmediaimage] and So-Fake-R1 [huang2025so] unify tampered and synthetic detection in a multi-class setting to evaluate MLLM capabilities. Collectively, these efforts mark a shift in IFD research toward multimodal reasoning, yielding more accurate and interpretable forgery analysis.
@@ 2.2 Thinking with Images
The “Thinking with Images” paradigm is pushing MLLMs beyond passive description toward interactive, iterative agents [zhu2024intelligent]. Early work typically relied on predefined CoT formats and static toolsets (e.g., VisProg [gupta2023visual] and ViperGPT [suris2023vipergpt]) to prompt models to call fixed tools for specific vision tasks, but this design limits flexibility and generalization. To address this, METATOOL [wang2024metatool] introduces meta-task augmentation to improve tool mastery and transfer, while PyVision [zhao2025pyvision] enables dynamic Python code generation to invoke complex tools for more versatile reasoning. Pushing autonomy further, recent studies optimize tool use via reinforcement learning (RL) [liu2025visual]. DeepEyes [zheng2025deepeyes] performs end-to-end reinforcement learning with tool-oriented data selection and reward design; V-TOOLRL [su2025openthinkimg] directly maximizes task success from tool-interaction feedback; and ReVPT [zhou2025reinforced] adopts a two-stage scheme with multiple visualization tools to lift general capability.
@@ 3.1 FABench
Recent advances in generative AI have enabled the easy creation of sophisticated synthetic and tampered content, while existing detection datasets exhibit notable limitations: (i) Outdated synthetic content. Benchmarks grounded in early GANs (e.g., StyleGAN [karras2019style]) mainly contain low-fidelity generations that are significantly easier than modern photorealistic outputs (e.g., GPT-4o-image [jaech2024openai], Midjourney-v7 [midjourney2023]). (ii) Fixed tampering pipelines. Many datasets rely on a narrow set of inpainting models (e.g., Stable Diffusion [sdxl]) and rarely explore newer pipelines (e.g., FLUX-Kontext [flux2024], Qwen-image [wu2025qwenimage]), limiting heterogeneity and inducing repeated artifacts. We build FABench via a strict, modular pipeline designed to maximize diversity across contemporary generators. FABench contains authentic, synthetic, and tampered images to reflect open-world scenarios and comprehensively evaluate forensic reasoning:
@@
  • Tampered (30k): Starting from COCO sources with instance masks, we derive object masks via SAM-guided text prompts [lin2014microsoft], then perform object-level inpainting using (i) strict-mask models (FLUX-1-Fill, Stable Diffusion; inputs: image/mask/prompt) and (ii) soft/no-mask models (e.g., GPT-4o-image, Qwen-Image), where the edited regions are composited back into the original image to suppress unintended global micro-changes. The pipeline is provided in the Supplementary Material.
We adopt a multi-stage pipeline consisting of quality validation (resolution bounds, file integrity, mask legality, etc.) and deduplication, followed by stratified human auditing to filter low-quality samples. Samples that fail inspection are removed, while borderline cases are re-synthesized or re-inpainted accordingly. For the tampered split, we construct a 700-image tampered test set using seven generators: GPT-4o, DALL·E 3 [openai2024dalle3], FLUX-1-dev, FLUX-Kontext, Stable Diffusion, Qwen-Image, and Nano Banana [nano]. Similarly, for the synthetic split, we generate a 700-image synthetic test set using GPT-4o, DALL·E 3, FLUX-1-dev, Midjourney-v7, Stable Diffusion, Qwen-Image, and Nano Banana. For the authentic split, we randomly sample 700 real images from COCO to form the authentic test set. The composition of the training set is provided in the Supplementary Material. Figure 3 further showcases examples of tampered and synthetic images produced by different generators in FABench. We observe that advanced models, such as Nano Banana and Qwen-Image, produce more photorealistic and harder-to-discriminate forgeries.
@@ 3.2 ForenAgent
We present ForenAgent, an novel interactive multi-turn framework that enables the MLLMs to autonomously generate, execute, and iteratively refine Python-based low-level tools for IFD. This approach provides more flexible and interpretable solutions compared to traditional methods. ForenAgent empowers MLLMs to dynamically generate and execute low-level Python code during the reasoning process. In each session, the MLLM receives input, generates Python code as a response, and executes it within an isolated Python runtime environment. The generated outputs, whether textual, visual, or both, are fed back into the MLLM’s context to iteratively refine its reasoning over multiple turns until a final answer is produced.
@@ 3.2.1 Tool Boxes
ForenAgent provides Python as the fundamental building block for tool construction. We identify two major categories of tools in the following:
(1) Basic Image Processing: These tools form the basis for visual manipulation and perception. They enable the agent to clean, align, and highlight image content to improve downstream reasoning.
@@
  • Enhancement: In visually subtle domains like tampered imaging, the agent applies contrast adjustments and other enhancements to make latent structures more prominent.
@@
  • Frequency Domain Analysis: Tools that analyze artifacts in transformed domains. (1) FFT High-Frequency Residual: Emphasizes forgery boundaries and texture anomalies in the frequency domain. (2) DWT High-Frequency Subbands: Uses wavelet decomposition to reveal high-frequency differences from synthesis or upsampling. (3) Resampling Periodicity: Detects spectral peaks introduced by interpolation (scaling/rotation). (4) DCT-based High-Pass Filter: Extracts high-frequency components to highlight edges and tampering traces.
  • Noise & Residual Analysis: Tools that extract subtle noise patterns typically suppressed by image content. (5) SRM: Uses a bank of high-pass and directional filters to extract robust noise residuals. (6) Bayar Constrained Convolution: Employs a specific convolutional kernel to suppress image content and amplify manipulation traces. (7) PRNU (Photo-Response Non-Uniformity): Extracts the camera sensor’s unique fingerprint noise to find local inconsistencies (splices) via block correlation.
  • Edge & Boundary Analysis: Methods to pinpoint inconsistent edges or gradients. (8) Sobel Edge Detector: Identifies splicing boundaries or anomalous edge patterns. (9) General High-Pass Filters: Extracts high-frequency components to detect tampering artifacts.
  • Specific Artifact Detection: Tools targeting the byproducts of distinct manipulations. (10) JPEG Ghost: Detects recompression artifacts by analyzing the error layer difference between multiple compression qualities. (11) Median Filtering Traces: Statistically measures artifacts and suspicious smoothing patterns.
  • Statistical Analysis: (12) Local Correlation Map: Quantifies enhanced correlations within pixel neighborhoods, often indicative of manipulation.

### v2

@@ Abstract
Existing image forgery detection (IFD) methods either exploit low-level, semantics-agnostic artifacts or rely on multimodal large language models (MLLMs) with high-level semantic knowledge. Although naturally complementary, these two information streams are highly heterogeneous in both paradigm and reasoning, making it difficult for existing methods to unify them or effectively model their cross-level interactions. To address this gap, we propose ForenAgent, a multi-round interactive IFD framework that enables MLLMs to autonomously invoke, execute, and iteratively refine Python-based low-level tools around the detection objective, thereby achieving more flexible and interpretable forgery analysis. ForenAgent adopts a two-stage training pipeline with Cold Start and Reinforcement Fine-Tuning to progressively improve tool interaction and reasoning adaptability. We design a human-inspired dynamic reasoning loop with global perception, local focusing, iterative probing, and holistic adjudication, and implement it as both a data-sampling strategy and a task-aligned process reward. To support training and evaluation, we built FABench, a heterogeneous agent-forensics dataset with 100k images and about 200k agent-interaction question-answer pairs. Experiments show that ForenAgent exhibits emergent tool-use competence and reflective reasoning on challenging IFD tasks when assisted by low-level tools, charting a promising route toward general-purpose IFD. The code is available at https://github.com/zfr00/ForenAgent.
@@ 1 Introduction
In this work, we propose ForenAgent, an interactive multi-turn framework that enables MLLMs to autonomously invoke, execute, and iteratively refine Python-based low-level tools for IFD. To this end, we abstract and generalize common low-level forensics operators, including frequency residual, noise residual, and high-pass filtering, and package them into a toolbox of 12 utilities for future extension. For efficiency, only basic image processing operations, such as cropping, require code generation, while the remaining low-level forensics tools are exposed via direct tool calls. As illustrated in Figure 1, ForenAgent autonomously orchestrates Python tools to verify a forged image from global screening to local inspection, ultimately classifying it as tampered and accurately localizing the forged region. The agent further demonstrates reflective self-correction by recovering from an initially misfocused crop to the correct area of interest, an “aha moment” observed in IFD agents.
The development of ForenAgent involves two key components: (1) Forgery Agent Benchmark (FABench), a high-quality and heterogeneous forensic agent dataset constructed using state-of-the-art generative models (e.g., GPT-4o [achiam2023gpt], Nano-Banana [nano], and Midjourney-v7 [midjourney2023])). It contains 100k images (40k real, 30k synthetic, and 30k tampered) and serves as a comprehensive benchmark for training and evaluation in IFD. (2) A Cold-Start and Reinforcement Fine-Tuning (RFT) framework, designed to train MLLMs to function as reliable and autonomous agents. During the Cold-Start stage, ForenAgent adopts a self-exploration and experience-distillation paradigm. Specifically, GPT-4.1 [openai2025gpt41] observes a large collection of forgery samples from FABench under system prompts that provide procedural guidance and executable code examples, distilling operational patterns into structured agent–interaction training data for initialization. During RFT, we abstract the human IFD workflow into four reasoning stages: global perception, local focusing, iterative probing, and holistic adjudication. Correspondingly, we design four Forgery Process Rewards that together form the overall tool reward. By incorporating these verifiable reward components into the reinforcement learning process, ForenAgent develops a more interpretable and systematic forensic reasoning mechanism, effectively integrating basic image processing with low-level forensic analysis in a coherent investigative workflow. This design enables ForenAgent to explore diverse reasoning strategies and optimize for long-term process quality rather than simply imitating predefined answers.
@@
(1) We propose ForenAgent, a novel interactive, multi-turn framework that enables an MLLM to autonomously invoke, execute, and iteratively refine Python-based low-level tools for image forgery detection, thereby taking the first step toward intelligent, tool-augmented IFD systems. (2) We construct FABench, a large-scale, high-quality, and heterogeneous forensic agent dataset comprising 100k images and 200k interactive QA pairs, focusing on forgery detection from cutting-edge generative models. (3) We formulate a dynamic reasoning loop comprising global perception, local focusing, iterative probing, and holistic adjudication into a data-sampling strategy and task-aligned process reward that foster flexible, tool-adaptive evidence reasoning and enable robust, interpretable decisions.
@@ 2.1 Image Forgery Detection
Early IFD research primarily relies on low-level feature extractors that model non-semantic inconsistencies between manipulated and authentic regions. Typical frequency and residual cues include FFT-based high-frequency modeling in FreqNet [tan2024frequency], SRM-filtered noise residuals in RGB-N [luo2021generalizing], multi-scale high-frequency noise suppression in HFF [luo2021generalizing], and transformation-driven feature learning in IT-Detector [li2025improving]. Related signals further exploit camera fingerprints (PRNU) [zhang2023prnu], inconsistent compression (JPEG ghost) [popescu2005exposing], DCT high-pass localization (ObjectFormer) [Objectformerwang2022objectformer], and boundary-sensitive operators (MVSS-Net, HiFi-Net) [MVSS-Net2021image, HiFi_IFDL]. Additional work broadens these cues and improves robustness [chen2022self, bayar2016deep, liu2024forgery]. ST-Trace analyzes images through spatial-temporal evolution patterns [wang2025spatial]. Despite progress, generalization remains challenging: Chameleon highlights failures on high-quality synthetic images [yan2024sanity], while Effort [yan2024orthogonal] improves robustness by separating content from artifact features, better isolating forgery traces. More recently, vision-language models and LLM-augmented frameworks have been introduced to incorporate semantic reasoning and natural-language explanations. DD-VQA [zhang2024common] enables fine-tuning BLIP on tampering data to improve both detection and explanations; FakeShield [xu2024fakeshield] and ForgeryGPT [liu2024forgerygpt] leverage LLMs for multimodal understanding and interactive, explainable analysis. For synthetic image detection, FakeScope [li2025fakescope] and LEGION [kang2025legion] further demonstrate the promise of MLLMs, while SIDA [huang2025sidasocialmediaimage] and So-Fake-R1 [huang2025so] unify tampered and synthetic detection in a multi-class setting to probe MLLM capabilities. Overall, the field is moving from artifact-centric detection toward multimodal reasoning with improved interpretability. AIGI-Holmes [zhou2025aigi] employs MLLMs for explainable AI-generated image detection, unifying high-level reasoning with fine-grained artifact analysis to boost generalization. Loki [ye2024loki] establishes a comprehensive benchmark for synthetic data detection, exposing the disparity between MLLMs’ high-level reasoning and fine-grained artifact perception.
@@ 2.2 Thinking with Images
The “Thinking with Images” paradigm is pushing MLLMs beyond passive description toward interactive, iterative agents [zhu2024intelligent]. Early work typically relied on predefined CoT formats and static toolsets (e.g., VisProg [gupta2023visual] and ViperGPT [suris2023vipergpt]) to prompt models to call fixed tools for specific vision tasks, but this design limits flexibility and generalization. To address this, METATOOL [wang2024metatool] introduces meta-task augmentation to improve tool mastery and transfer. In contrast, PyVision [zhao2025pyvision] enables dynamic Python code generation to invoke complex tools for more versatile reasoning. Pushing autonomy further, recent studies optimize tool use via reinforcement learning (RL) [liu2025visual]. DeepEyes [zheng2025deepeyes] performs end-to-end reinforcement learning with tool-oriented data selection and reward design; DeepEyesV2 [hong2025deepeyesv2] evolves multimodal reasoning into an agentic paradigm, achieving SOTA results on perception and search-intensive tasks within the RealX-Bench. Thyme [zhang2025thyme] shifts focus from static image perception to comprehensive reasoning by integrating temporal and logical context. V-TOOLRL [su2025openthinkimg] directly maximizes task success from tool-interaction feedback, and ReVPT [zhou2025reinforced] adopts a two-stage scheme with multiple visualization tools to lift general capability.
@@ 3.1 FABench
Recent advances in generative AI have enabled the easy creation of sophisticated synthetic and tampered content, while existing detection datasets exhibit notable limitations: (i) Outdated synthetic content. Benchmarks grounded in early GANs (e.g., StyleGAN [karras2019style]) mainly contain low-fidelity generations that are significantly easier than modern photorealistic outputs (e.g., GPT-4o-image [achiam2023gpt], Midjourney-v7 [midjourney2023]). (ii) Fixed tampering pipelines. Many datasets rely on a narrow set of inpainting models (e.g., Stable Diffusion [sdxl]) and rarely explore newer pipelines (e.g., FLUX-Kontext [flux2024], Qwen-image [wu2025qwenimage]), limiting heterogeneity and inducing repeated artifacts. We build FABench via a strict, modular pipeline designed to maximize diversity across contemporary generators. FABench contains authentic, synthetic, and tampered images to reflect open-world scenarios and comprehensively evaluate forensic reasoning. Beyond simple objects and portraits, it also covers a wide range of naturalistic scenes:
@@
Tampered (30k): Starting from COCO sources with instance masks, we derive object masks via SAM-guided text prompts [lin2014microsoft], then perform object-level inpainting using (i) strict-mask models (FLUX-1-Fill, Stable Diffusion; inputs: image/mask/prompt) and (ii) soft/no-mask models (e.g., GPT-4o-image, Qwen-Image), where the edited regions are composited back into the original image to suppress unintended global micro-changes. The comprehensive construction pipeline and detailed analysis are provided in the Appendix.
We adopt a multi-stage pipeline consisting of quality validation (file integrity, mask legality, etc.) and deduplication, followed by stratified human auditing to filter low-quality samples. Samples that fail inspection are removed, while borderline cases are re-synthesized or re-inpainted accordingly. For the tampered split, we construct a 700-image tampered test set using seven generators: GPT-4o, DALL·E 3 [openai2024dalle3], FLUX-1-Fill, FLUX-Kontext, Stable Diffusion, Qwen-Image, and Nano Banana [nano]. Similarly, for the synthetic split, we generate a 700-image synthetic test set using GPT-4o, DALL·E 3, FLUX-1-dev, Midjourney-v7, Stable Diffusion, Qwen-Image, and Nano Banana. For the authentic split, we randomly sample 700 real images from COCO to form the authentic test set.
We visualize the composition of the FABench training set in Figure 3. For synthetic images, the dataset includes samples generated by GPT-4o (4 k), DALL·E 3 (2 k), FLUX-1-dev (5 k), Midjourney-v7 (5 k), Stable Diffusion (5 k), Qwen-Image (5 k), and Nano Banana (4 k). For tampered images, it includes GPT-4o (2 k), DALL·E 3 (2 k), FLUX-1-Fill (10 k), FLUX-Kontext (2 k), Stable Diffusion (10 k), Qwen-Image (2 k), and Nano Banana (2 k). This diverse composition demonstrates that FABench integrates a wide spectrum of generation and inpainting paradigms, covering both text-to-image and mask-based pipelines. Figure 4 further showcases examples of tampered and synthetic images produced by different generators in FABench. We observe that advanced models, such as Nano Banana and Qwen-Image, produce more photorealistic and harder-to-discriminate forgeries.
@@ 3.2 ForenAgent
ForenAgent enables MLLMs to dynamically invoke and execute low-level tools throughout their reasoning process. In each session, the MLLM receives the input and responds with either executable Python code or structured tool calls, which are run in an isolated Python sandbox. The resulting outputs, including text and or visualizations, are then fed back into the MLLM context, allowing iterative multi-turn refinement until the final prediction is produced.
@@ 3.2.1 Tool Boxes
ForenAgent adopts Python code and tool calls as the fundamental primitives for tool construction. Accordingly, we categorize the tools into two major types as follows:
(1) Basic Image Processing: These two categories of tools form the basis for visual manipulation and perception. They enable the agent to clean, align, and highlight image content to improve downstream reasoning.
@@
Enhancement: In visually subtle domains like tampered imaging, the agent autonomously applies adaptive contrast adjustments and multi-scale enhancements to amplify latent structures and imperceptible forensic traces, making them significantly more prominent.
@@
(a) Frequency Domain Analysis: Tools that analyze artifacts in transformed domains. (1) FFT High-Frequency Residual: Emphasizes forgery boundaries and texture anomalies in the frequency domain. (2) DWT High-Frequency Subbands: Uses wavelet decomposition to reveal high-frequency differences from synthesis or upsampling. (3) Resampling Periodicity: Detects spectral peaks introduced by interpolation (scaling/rotation). (4) DCT-based High-Pass Filter: Extracts high-frequency components to highlight edges and tampering traces.
(b) Noise & Residual Analysis: Tools that extract subtle noise patterns typically suppressed by image content. (5) SRM: Uses a bank of high-pass and directional filters to extract robust noise residuals. (6) Bayar Constrained Convolution: Employs a specific convolutional kernel to suppress image content and amplify manipulation traces. (7) PRNU (Photo-Response Non-Uniformity): Extracts the camera sensor’s unique fingerprint noise to find local inconsistencies (splices) via block correlation.
(c) Edge & Boundary Analysis: Methods to pinpoint inconsistent edges or gradients. (8) Sobel Edge Detector: Identifies splicing boundaries or anomalous edge patterns. (9) Laplacian High-Pass: Extracts high-frequency components to detect tampering artifacts.
(d) Specific Artifact Detection: Tools targeting the byproducts of distinct manipulations. (10) JPEG Ghost: Detects recompression artifacts by analyzing the error layer difference between multiple compression qualities. (11) Median Filtering Traces: measures artifacts and suspicious smoothing patterns.
(e) Statistical Analysis: (12) Local Correlation Map: Quantifies enhanced correlations within pixel neighborhoods, often indicative of manipulation.
For efficiency, only basic image processing operations such as cropping require Python code generation and execution, while the remaining low level forensics tools are exposed via direct tool calls.